A body 'A' is going from south to north and body 'B' is going west to east. Determine direction of relative velocity of 'A' with respect to 'B'.

Grade:11

4 Answers

vibhu mittal
13 years ago

north west.

vijesh jayan
13 years ago
upward
deepshikha k
12 years ago

upward as dir. of relative velocity ofa wrt b=90-0=90

deepshikha k
12 years ago

upward as dir. of rel vel. of a w.r.t. b =90-0=90